# Large Language Models as Virtual Survey Respondents: Evaluating Sociodemographic Response Generation

### v1

@@ Abstract.
Questionnaire-based surveys are foundational to social science research and public policymaking, yet traditional survey methods remain costly, time-consuming, and often limited in scale. This paper explores a new paradigm: simulating virtual survey respondents using Large Language Models (LLMs). We introduce two novel simulation settings, namely Partial Attribute Simulation (PAS) and Full Attribute Simulation (FAS), to systematically evaluate the ability of LLMs to generate accurate and demographically coherent responses. In PAS, the model predicts missing attributes based on partial respondent profiles, whereas FAS involves generating complete synthetic datasets under both zero-context and context-enhanced conditions. We curate a comprehensive benchmark suite, LLM-S3 (Large Language Model-based Sociodemographic Survey Simulation), that spans 11 real-world public datasets across four sociological domains. Our evaluation of multiple mainstream LLMs (GPT-3.5/4 Turbo, LLaMA 3.0/3.1-8B) reveals consistent trends in prediction performance, highlights failure modes, and demonstrates how context and prompt design impact simulation fidelity. This work establishes a rigorous foundation for LLM-driven survey simulations, offering scalable and cost-effective tools for sociological research and policy evaluation. Our code and dataset are available at: https://github.com/dart-lab-research/LLM-S-Cube-Benchmark.
@@ 1. Introduction
Questionnaire-based surveys serve as critical instruments for understanding public sentiment, tracking social dynamics, and guiding evidence-based policymaking. However, the traditional survey pipeline faces mounting challenges: high operational costs, lengthy data collection timelines, and declining response rates (Mou et al., 2024). These limitations hinder the timely delivery of insights essential for responsive policy formulation. To address these issues, we propose and systematically investigate a novel approach: leveraging Large Language Models (LLMs) to simulate virtual survey respondents.
@@
Recent advances in LLMs have demonstrated remarkable generalization capabilities in tasks ranging from question answering to synthetic text generation (Brown et al., 2020; Grattafiori et al., 2024; Törnberg, 2023). Despite growing interest in LLM-driven simulations (Dillion et al., 2023; Kim and Lee, 2023; Jansen et al., 2023), existing research often suffers from three limitations that our work addresses: (1) narrow task coverage, where prior studies frequently focus on a limited number of survey items or demographic traits, restricting their generalizability across diverse tasks (Yang et al., 2024; Zhou et al., 2024; Sun et al., 2024; Argyle et al., 2022); (2) domain-specific constraints, as most existing works are tailored to particular sociological themes (e.g., politics or consumer behavior), which impedes comprehensive cross-domain evaluation (Chaudhary and Penn, 2024; Yu et al., 2024; Argyle et al., 2022; Törnberg et al., 2023; Cao et al., 2025); and (3) inadequate benchmarking, primarily due to reliance on synthetic or overly simplified datasets that fail to reflect the complexity or scale of real-world survey distributions (Argyle et al., 2022; Ziems et al., 2024; Ren et al., 2024; Cao et al., 2025). Our research bridges these gaps by introducing a systematic and extensible framework for evaluating LLMs in survey response generation.
To this end, we introduce two distinct and novel simulation paradigms: Partial Attribute Simulation (PAS), where LLMs predict missing demographic or behavioral attributes from incomplete respondent profiles, and Full Attribute Simulation (FAS), where LLMs generate synthetic datasets under varying degrees of contextual constraint. These paradigms are designed to reflect practical scenarios in applied research, from imputing nonresponse in existing surveys to simulating synthetic populations for complex policy experiments.
@@
To support reproducibility and broad applicability, we curate a cross-domain benchmark suite, LLM-S3 (Large Language Model-based Sociodemographic Survey Simulation), composed of eleven public datasets across four sociological areas: Social and Public Affairs, Work and Income, Household and Behavioral Patterns, and Health and Lifestyle. We evaluate four representative LLMs (GPT-3.5 Turbo, GPT-4 Turbo, LLaMA 3.0-8B, LLaMA 3.1-8B) on both numerical and categorical prediction tasks under zero-shot and few-shot settings. Our extensive experiments yield several key findings: (i) LLMs exhibit reliable and consistent performance trends across various model families and task types; (ii) prompt design and contextual augmentation significantly impact simulation fidelity, highlighting their crucial role; and (iii) failure cases in structured output generation remain a major bottleneck, particularly evident in the Full Attribute Simulation (FAS) scenario. These insights not only advance our fundamental understanding of LLMs in survey applications but also provide practical design principles for future research and deployment in this emerging field.
In sum, to the best of our knowledge, this work offers the first large-scale, systematic exploration of LLMs as virtual survey respondents, contributing both a novel simulation methodology and a robust, multi-domain benchmark to drive future advances in LLM-assisted social research.
@@ 2. Related Work
2.1. LLM Role-Playing and Personalization
@@
Recent advancements in LLMs have improved the realism and contextual relevance of agent behaviors across domains. A key focus has been on persona-driven role-playing, where maintaining coherent persona traits and consistent decision-making is essential (Peng and Shang, 2024; Wang et al., 2024; Xu et al., 2024). To enhance personalization further, user-specific information such as traits and profiles has been incorporated into dialogue agents, enabling responses that better reflect individual users (Zhang et al., 2018; Salemi et al., 2024). However, integrating such personalized data with large language models remains challenging, which highlights ongoing opportunities for improving adaptability (Chen et al., 2024). Complementary efforts have developed evaluation frameworks such as ECHO to measure how well LLMs replicate human-like personalized responses, thereby advancing the authenticity and naturalness of agent interactions (Ng et al., 2024; Lotfi et al., 2024).
2.2. Using LLMs for Sociological Simulation
LLMs have shown promise in simulating complex social behaviors and demographic patterns, especially in public opinion dynamics and demographic predictions. For example, prior work has simulated large-scale opinion dynamics through modeling interactions among over 1,000 virtual citizens (Park et al., 2024) and applied multi-agent reasoning for election predictions (Yu et al., 2024). In parallel, other methods employ demographic-conditioned sampling to simulate subgroup opinions; however, validation against longitudinal survey data remains limited (Argyle et al., 2022). Some studies focus on simulating group-level statistics using conditional sampling techniques, while others predict individual-level attributes such as emotions or personality traits based on psychological metrics (Huang et al., 2023; Pan and Zeng, 2023; Xie et al., 2025). Applications also include inferring psychological dispositions from textual traces and predicting social attributes through prompt engineering (Chaudhary and Penn, 2024; Zhou et al., 2024; Egami et al., 2024; Rosenbusch et al., 2023).
@@ 3. Partial Attribute Simulation (PAS)
In numerous real-world survey scenarios, researchers frequently possess foundational demographic information about respondents (e.g., age, gender, occupation) but lack specific insights into individual opinions or preferences on particular topics. This inherent data sparsity poses a significant challenge for accurately simulating survey responses under such constraints. To address this, we introduce the Partial Attribute Simulation (PAS) setting, a novel framework designed to evaluate LLMs’ ability to predict missing sociodemographic attributes. As illustrated in Figure 1, PAS operates by providing LLMs with only fundamental demographic data and selected conditional attributes, without access to an individual’s complete historical response profile. The LLMs are then tasked with generating plausible responses, effectively simulating how individuals with specific partial characteristics might answer survey questions. This paradigm offers valuable insights into the influence of partial information on survey responses, enabling the simulation of realistic survey scenarios across diverse sociological contexts.
@@ 3.2. Evaluation Strategies in PAS
[…]
@@ 4. Full Attribute Simulation (FAS)
Large-scale sociological analysis frequently necessitates understanding population-level trends and aggregate distributions, beyond individual-level responses. While the PAS scenario predicts specific attributes for individual agents, the Full Attribute Simulation (FAS) tackles a more complex challenge: generating comprehensive synthetic datasets that accurately preserve demographic distributions while inferring unobserved attributes. As illustrated in Figure 1, this capability is critical for policy formulation when only aggregated statistics (e.g., age/gender distributions) are available, but granular, domain-specific insights (e.g., opinions on urban development) remain unobserved. FAS presents a compelling paradigm for synthesizing complete, statistically representative populations, offering a robust foundation for sophisticated analytical tasks in social science research and evidence-based policy design.

### v2

@@ Abstract.
Questionnaire-based surveys are foundational to social science research and public policymaking, yet traditional survey methods remain costly, time-consuming, and often limited in scale. Although prior work has explored large language models (LLMs) as virtual survey respondents, existing studies often address narrow task settings, focus on single sociological domains, or lack a unified evaluation framework that enables systematic comparison across diverse datasets and models. To address these gaps, we introduce two complementary task abstractions: Partial Attribute Simulation (PAS), where LLMs predict missing attributes from incomplete respondent profiles, and Full Attribute Simulation (FAS), where LLMs generate complete synthetic datasets under zero-context and context-enhanced conditions. Both are framed as diagnostic and exploratory tools rather than replacements for human data collection. We curate LLM-S3 (Large Language Model-based Sociodemographic Survey Simulation), a benchmark spanning 11 real-world public datasets across four sociological domains, and evaluate GPT-3.5/4 Turbo and LLaMA 3.0/3.1-8B under zero-shot and few-shot settings. Our evaluation reveals consistent performance trends across model families, highlights failure modes in structured output generation, and demonstrates how context and prompt design affect simulation fidelity. Our code and dataset are available at: https://github.com/dart-lab-research/LLM-S-Cube-Benchmark.
@@ 1. Introduction
Questionnaire-based surveys serve as critical instruments for understanding public sentiment, tracking social dynamics, and guiding evidence-based policymaking. However, the traditional survey pipeline faces mounting challenges: high operational costs, lengthy data collection timelines, and declining response rates (Mou et al., 2024). These limitations hinder the timely delivery of insights essential for responsive policy formulation. To address these issues, we investigate a structured framework for leveraging Large Language Models (LLMs) in survey response generation.
@@
Recent advances in LLMs have demonstrated remarkable generalization capabilities in tasks ranging from question answering to synthetic text generation (Brown et al., 2020; Grattafiori et al., 2024; Törnberg, 2023). While there is growing interest in LLM-driven simulations (Dillion et al., 2023; Kim and Lee, 2023; Jansen et al., 2023) and prior work has explored LLMs as survey respondents (Argyle et al., 2023; Park et al., 2024; Bisbee et al., 2024), existing research often suffers from three limitations that our work addresses: (1) narrow task coverage, where prior studies frequently focus on a limited number of survey items or demographic traits, restricting their generalizability across diverse tasks (Yang et al., 2024; Zhou et al., 2024; Sun et al., 2024; Argyle et al., 2023); (2) domain-specific constraints, as most existing works are tailored to particular sociological themes (e.g., politics or consumer behavior), which impedes comprehensive cross-domain evaluation (Chaudhary and Penn, 2024; Yu et al., 2024; Argyle et al., 2023; Törnberg et al., 2023; Cao et al., 2025); and (3) inadequate benchmarking, primarily due to reliance on synthetic or overly simplified datasets that fail to reflect the complexity or scale of real-world survey distributions (Argyle et al., 2023; Ziems et al., 2024; Ren et al., 2024; Cao et al., 2025). Our research bridges these gaps by introducing a systematic and extensible framework for evaluating LLMs in survey response generation.
To this end, we introduce two distinct and complementary task settings: Partial Attribute Simulation (PAS), where LLMs predict missing demographic or behavioral attributes from incomplete respondent profiles, and Full Attribute Simulation (FAS), where LLMs generate synthetic datasets under varying degrees of contextual constraint. These paradigms are designed to reflect practical scenarios in applied research. For instance, PAS addresses settings where a government agency possesses detailed demographic records (e.g., age, income, occupation) for its residents but lacks information on specific opinions or preferences, such as attitudes toward public transportation or renewable energy. By conditioning on the known attributes, PAS can generate plausible predictions for the missing survey responses, enabling population-level trend analysis without conducting a full-scale survey. In contrast, FAS targets scenarios where an organization, such as a market research firm, has access to aggregate statistics about a target population (e.g., age distributions and income brackets) but cannot obtain individual-level profiles. FAS can then synthesize a statistically representative synthetic population, supporting exploratory analyses such as consumer preference estimation or policy impact assessment. We clarify that PAS and FAS are best understood as diagnostic, methodological, or exploratory tools rather than a viable replacement for human survey data collection.
@@
To support reproducibility and broad applicability, we curate a cross-domain benchmark suite, LLM-S3 (Large Language Model-based Sociodemographic Survey Simulation), composed of eleven public datasets across four sociological areas: Social and Public Affairs, Work and Income, Household and Behavioral Patterns, and Health and Lifestyle. We note that the datasets used in this study are predominantly drawn from North American and European contexts, reflecting the current availability of large-scale public survey data. While this scope is a practical constraint of the present evaluation rather than a fundamental limitation of the PAS/FAS framework itself, the benchmark is designed to be extensible: as cross-cultural and multilingual survey datasets become available, the same evaluation protocol can be applied to assess LLM performance across diverse sociological contexts. We evaluate four representative LLMs (GPT-3.5 Turbo, GPT-4 Turbo, LLaMA 3.0-8B, LLaMA 3.1-8B) on both numerical and categorical prediction tasks under zero-shot and few-shot settings. Our extensive experiments yield several key findings: (i) LLMs exhibit reliable and consistent performance trends across various model families and task types; (ii) prompt design and contextual augmentation significantly impact simulation fidelity, highlighting their crucial role; and (iii) failure cases in structured output generation remain a major bottleneck, particularly evident in the Full Attribute Simulation (FAS) scenario. We explicitly acknowledge that our evaluation measures statistical and distributional similarity, not behavioral fidelity, construct validity, or the nuanced reasoning of real human respondents. These insights not only advance our fundamental understanding of LLMs in survey applications but also provide practical design principles for future research and deployment in this emerging field.
@@ 2. Related Work
2.1. From Role-Playing to Structured Survey Simulation
@@
Recent advancements in LLMs have enabled increasingly realistic agent behaviors through persona-driven role-playing, where maintaining coherent persona traits and consistent decision-making is essential (Peng and Shang, 2024; Wang et al., 2024; Xu et al., 2024). User-specific information such as traits and profiles has been incorporated into dialogue agents to enhance personalization (Zhang et al., 2018; Salemi et al., 2024), and evaluation frameworks have been developed to measure how well LLMs replicate human-like personalized responses (Ng et al., 2024; Lotfi et al., 2024). These efforts demonstrate that LLMs can effectively simulate human behavior when grounded in appropriate contextual and demographic information.
The natural extension of this capability is to apply LLM role-playing to sociological research, using models as virtual survey respondents. However, a critical mismatch emerges between the two domains. Role-playing and persona-faithfulness studies (Peng and Shang, 2024; Wang et al., 2024; Xu et al., 2024) are designed for open-domain dialogue or narrative settings, where the goal is to maintain character consistency over extended interactions. Survey simulation, by contrast, requires structured, attribute-conditioned prediction: the model must infer a specific Likert-scale response given a partial demographic profile, without the benefit of conversational context or extended character development. The metrics that work for dialogue consistency (persona faithfulness, interview fidelity) do not transfer to the task of predicting survey responses from structured demographic inputs. This fundamental mismatch motivates the need for a distinct framework that formalizes survey simulation as a rigorous, reproducible task with its own evaluation criteria.
2.2. Survey Simulation: From Feasibility to Methodological Rigor
Research on LLM-based survey simulation has progressed through three distinct phases, each revealing both capabilities and critical limitations that shape our understanding of what is required for reliable simulation.
Phase 1: Establishing Feasibility. Early work demonstrated that LLMs can be prompted to reproduce broad attitudinal patterns across demographic subgroups. Argyle et al. (Argyle et al., 2023) showed that models encode sufficient sociological knowledge to simulate survey responses, establishing the “silicon sampling” paradigm. While this established basic feasibility, it also raised immediate questions: What are the conditions under which LLM responses reliably track human distributions? What factors limit generalizability?
Phase 2: Pushing Toward Behavioral Realism. Subsequent work addressed these questions by grounding simulations in richer individual-level data. Park et al. (Park et al., 2024) constructed detailed agent memories from interview data, enabling virtual respondents to exhibit behavioral consistency over time. This approach demonstrated that individual-level grounding improves simulation quality, but the requirement of extensive per-individual data collection limits scalability and practical applicability.
Phase 3: Evaluating Validity. The most recent work has introduced methodological rigor by explicitly testing whether LLM-generated data can substitute for human responses in actual research workflows. Bisbee et al. (Bisbee et al., 2024) found that even when distributions appear statistically similar, they can lead to substantively different inferential conclusions. This finding is pivotal: it establishes that distributional similarity is necessary but not sufficient for valid survey simulation, and that the gap between statistical fidelity and behavioral fidelity is an open methodological challenge requiring systematic investigation.
Summary: Persistent Structural Gaps. While these three phases have advanced our understanding of LLM capabilities in survey simulation, the field has not converged on a shared conceptual framework or evaluation standard. This fragmentation leads to three specific structural gaps: (1) lack of conceptual clarity about what tasks survey simulation should perform (predicting missing attributes from partial profiles versus generating full synthetic populations from aggregate statistics); (2) no multi-domain benchmark enabling consistent evaluation across diverse sociological settings; and (3) no standardized evaluation protocols, hindering fair comparison across studies.
Our work addresses these gaps directly. The PAS/FAS abstraction brings conceptual clarity to two distinct simulation regimes: partial versus full attribute simulation. The LLM-S3 benchmark provides a unified, multi-domain evaluation suite spanning eleven real-world datasets across four sociological domains, enabling systematic cross-domain comparison that prior work has not supported.
@@ 3. Partial Attribute Simulation (PAS)
In numerous real-world survey scenarios, an organization already holds structured demographic records for a population of interest but lacks the specific attitudinal or behavioral measurements it needs. For example, a government labor agency might maintain administrative records for thousands of employees, including age, gender, job level, education, and employment type, yet lack survey data on job satisfaction, stress, or work-life balance. Conducting a full survey is costly and time-consuming; however, the agency still needs distributional estimates of these outcomes to inform workforce policy. PAS is designed for precisely this setting: the available demographic attributes serve as the input profile, while the unmeasured attitudinal traits act as the prediction targets. Conditioned on each individual’s known profile, the LLM is prompted to predict the missing responses, enabling population-level trend analysis without a full-scale survey. Crucially, PAS does not require any individual-level response history. It relies solely on the demographic data that the organization already possesses, which distinguishes it from collaborative filtering or imputation methods that require at least partial response data for the target attributes. As illustrated in Figure 1, PAS operates by providing LLMs with only fundamental demographic data and selected conditional attributes, without access to an individual’s complete historical response profile. The LLMs are then tasked with generating plausible responses, effectively simulating how individuals with specific partial characteristics might answer survey questions. This paradigm offers valuable insights into the influence of partial information on survey responses, enabling the simulation of realistic survey scenarios across diverse sociological contexts.
@@ 3.2. Evaluation Strategies in PAS
| ANES | 0.78540.0598 | 0.83580.0324 | 0.78270.0487 | 0.75250.0372 | 0.77960.0447 | 0.82190.0395 | 0.76860.0382 | 0.73550.0491 |
| GSS | 0.55880.1359 | 0.63130.0340 | 0.52890.0583 | 0.49690.0326 | 0.49480.0430 | 0.62230.0349 | 0.49330.0895 | 0.41290.0778 |
| BIS | 0.18650.0032 | 0.21650.0022 | 0.26290.0600 | 0.28480.0217 | 0.14650.0164 | 0.21690.0045 | 0.22610.0233 | 0.20770.0044 |
| RECS | 0.62080.1116 | 0.85980.0793 | 0.72570.0144 | 0.65400.0730 | 0.63860.0257 | 0.91130.0165 | 0.74830.2132 | 0.67420.0623 |
| EmpS | 0.21910.0293 | 0.21710.0017 | 0.23020.0259 | 0.21880.0183 | 0.22670.0092 | 0.21160.0105 | 0.21470.0230 | 0.21430.0045 |
| NHTS | 0.73140.0077 | 0.72310.0003 | 0.57320.0126 | 0.65070.0453 | 0.72460.0322 | 0.73000.0057 | 0.64090.0797 | 0.59870.0270 |
| Trell SMU | 0.64150.0105 | 0.65250.0050 | 0.38650.0138 | 0.50280.0134 | 0.72280.0357 | 0.66010.0057 | 0.70990.0051 | 0.73090.0176 |
| MxMH | 0.12830.0380 | 0.10210.0018 | 0.09010.0129 | 0.09330.0151 | 0.15420.0517 | 0.09770.0093 | 0.09200.0103 | 0.08600.0120 |
| YPS | 0.27290.0074 | 0.25590.0038 | 0.30830.0479 | 0.27870.0426 | 0.29160.0283 | 0.29680.0471 | 0.27680.0175 | 0.19900.0418 |
| MHD | 0.33770.0058 | 0.33750.0045 | 0.33900.0068 | 0.34340.0070 | 0.33140.0198 | 0.33390.0079 | 0.34030.0094 | 0.34370.0063 |
@@ 3.3. PAS Performance Analysis
3.4. Attribute Predictability and Demographic Profile Alignment in PAS
A central question for PAS is not merely whether an LLM can predict survey responses, but rather which attributes are predictable from a given demographic profile. Aggregate accuracy alone cannot answer this question, as it conflates attributes that are strongly determined by observable demographics with those governed by latent psychological or contextual factors. To surface this distinction, we conduct a subgroup-level analysis on the EmpS dataset () using LLaMA 3.1-8B in the zero-shot setting. We select EmpS because its five target attributes (WLB, WorkEnv, Workload, Stress, JobSatisfaction) all measure related workplace attitude constructs on a uniform 5-level ordinal scale, and its input profile includes rich, directly relevant demographic covariates (age, gender, job level, department, employment type, education), enabling controlled cross-attribute comparison where differences in predictability can be attributed to the target construct itself rather than to variation in response scales, domains, or input feature relevance. We use LLaMA 3.1-8B because, as a fully open-weight model, it permits local deployment, deterministic decoding, and reproducible repeated runs without dependence on proprietary API availability or versioning changes. We note that this analysis is a focused diagnostic case study rather than a comprehensive cross-model evaluation. Prediction accuracy is disaggregated by age group, gender, and job level across three independent runs (reported as meanstd). All five target attributes are 5-choice questions, yielding a random baseline accuracy of 20%.
Figure 3 presents the per-subgroup prediction accuracy. As reported in Table 2, the overall mean accuracy of LLaMA 3.1 on EmpS under zero-shot evaluation is , only marginally above the random baseline of . This result is itself informative: it reveals that the EmpS demographic profile (age, gender, job level, education, employment type) provides limited predictive signal for most workplace attitude attributes, suggesting that these outcomes are shaped primarily by individual experience and organizational context rather than by sociodemographic characteristics alone. Critically, however, the picture is not uniform across attributes. JobSatisfaction () consistently exceeds the baseline across all subgroups and demographic dimensions, indicating that this attribute maintains a detectable statistical association with the provided profile features. Stress (), by contrast, falls below the baseline, implying that stress levels are largely orthogonal to the demographic covariates available in EmpS. This attribute-level divergence constitutes a key insight enabled by the PAS benchmark: it allows practitioners to identify, prior to deployment, which survey constructs are amenable to demographic-profile-based simulation and which are not.
The subgroup breakdown further refines this picture. For Stress, the 51+ age group achieves only accuracy, a percentage-point gap below the 18–30 cohort (), and Lead-level employees are the hardest to predict (), falling percentage points below the Mid-level group (). These patterns suggest that stress responses among senior-age and leadership-role employees are particularly idiosyncratic, likely reflecting heterogeneous organizational pressures that demographic proxies cannot capture. For JobSatisfaction, the gender dimension reveals a percentage-point gap between the Other category () and Male respondents (), consistent with the hypothesis that satisfaction correlates more strongly with demographic identity for non-binary respondents in this dataset. Across all subgroups, the low variance across runs (std ) confirms that these patterns are stable and not artifacts of sampling noise.
Figure 4 compares the ground-truth and LLM-predicted response distributions for JobSatisfaction and Stress. Both attributes exhibit skewed empirical distributions: JobSatisfaction is right-skewed (concentrated at high satisfaction levels), whereas Stress is left-skewed (concentrated at low stress levels). Despite this structural similarity in the ground truth, the model’s performance diverges markedly between the two attributes. For JobSatisfaction, the predicted distribution closely tracks the empirical baseline, capturing the right-skewed shape with mass concentrated at levels 4 and 5. This distributional alignment confirms that JobSatisfaction maintains a detectable statistical association with the available demographic profile, making it a suitable target for PAS. For Stress, however, the model fails to reproduce the heavily left-skewed distribution, which is predominantly concentrated at level 1 (). Instead, the predicted distribution spreads mass more uniformly across levels 2 through 5, effectively inverting the dominant ground-truth category. This failure mode reveals that Stress levels are largely orthogonal to the demographic covariates available in EmpS and that the model cannot reliably infer this attribute from the provided profile features. The contrast between these two attributes underscores a fundamental insight for PAS deployment: the suitability of a target attribute for demographic-profile-based simulation depends critically on whether that attribute maintains a robust statistical association with the available covariates. Attributes like JobSatisfaction, which exhibit such associations, can be predicted with reasonable fidelity at both the individual and population levels. Conversely, attributes like Stress, which lack these associations, should be treated with caution in any downstream analysis relying on distributional fidelity.
@@ 4. Full Attribute Simulation (FAS)
While PAS assumes that individual demographic records are available, a different and more demanding scenario arises when an organization possesses only aggregate population statistics rather than individual-level data. Consider a market research firm that has access to published census summaries for a target region, such as the joint distribution of age, gender, and income brackets, but holds no individual survey records. The firm needs to generate a synthetic population of respondents whose collective attribute distribution matches the known aggregate statistics, so that downstream analyses such as consumer preference estimation or policy impact assessment can be conducted. This is the setting that FAS is designed to address: the LLM is given population-level priors (e.g., “43% of respondents are aged 25–44, median household income is $62,000”) and asked to generate a set of synthetic individuals whose joint attribute distribution approximates these priors. Large-scale sociological analysis frequently necessitates understanding population-level trends and aggregate distributions beyond individual-level responses, and FAS tackles precisely this challenge: generating comprehensive synthetic datasets that accurately preserve demographic distributions while inferring unobserved attributes. As illustrated in Figure 1, this capability is critical for policy formulation when only aggregated statistics (e.g., age/gender distributions) are available, but granular, domain-specific insights (e.g., opinions on urban development) remain unobserved. Unlike PAS, which conditions on individual records and predicts missing attributes per person, FAS requires only aggregate statistics and generates an entirely synthetic population from scratch. It is therefore most appropriate as a diagnostic, methodological, or exploratory tool, for example to test analytical pipelines, generate synthetic training data, or explore hypothetical population scenarios, rather than as a substitute for real survey data collection (Bisbee et al., 2024).